At the beginning of Year 1, the company buys box manufacturing equipment for $2,300. The estimated residual value is $100. The estimated useful life is 4 years. As shown below:

Straight line method

Straight line method Beg Book value Depreciation expense Accumulated depreciation Ending net book value
Yr 1 2300 (2300-100/4) = 550 550 1750
yr 2 1750 550 1100 1200
Yr 3 1200 550 1650 650
Yr 4 550 550 2200 100

Double decline method

Double decline method Beg Book value Depreciation expense Accumulated depreciation Ending net book value
Yr 1 2300 2300*50% = 1150 1150 1150
yr 2 1150 1150*50% = 575 1725 575
Yr 3 575 575*50% = 288 2013 287
Yr 4 287 187 2200 100

What Are the effects of deprecation methods above have on financial ratios shown below? Assume revenue is $5,000; operating income excluding deprecation is $3,00; and Property, Plant and Equipment (PPE) is the only total asset for the company.

Operating profit margin: Strait-line method. Accelerated method(Double Declining)

-Year 1

-Year 4

Return on Assets:

-Year 1

-Year 4

Asset turnover Ratio:   

-Year 1

-Year 4
Answer #1
Year-1 Year- 4 Analysis
St.line depn. DDB method depn. St.line depn. DDB method depn.
Operating Profit margin= Opg. Profit /Revenues
Revenues 5000 5000 5000 5000
Operating income excluding depn. 3000 3000 3000 3000
Less:Depn. 550 1150 550 187
Operating income 2450 1850 2450 2813
Operating Profit margin= Opg. Profit /Revenues 2450/5000= 1850/5000= 2450/5000= 2813/5000=
49% 37% 49% 56% Opg.Profit margin as per DDB reduced in initial yrs. & more in later yrs.compared to st.line depn.
Year- 4
Return on assets(ROA):
Opg. Income(after depn.)/Total assets
Total assets at end of year 1750 1150 100 100
Operating income 2450 1850 2450 2813
ROA 1.4 1.61 24.5 28.13 As net asset balance also decreases more than that under St.line method of depn.ROA is more under DDB than st.line, in both the years.
Asset turnover ratio(ATO):
Revenues/Assets
Revenues 5000 5000 5000 5000
Total assets at end of year 1750 1150 100 100
ATO 2.86 4.35 50 50 Due to more $ depn. under DDB, asset utilisation shows a better figure under the same, than under st.line method. When the asset has reached its salvage value, the asset utilisation, ie. $ revenue earned per $ of asset , is the same under both methods.